OpenStudy (anonymous):

(y + z)^3 I got y^3+z^3. Is that right?

OpenStudy (anonymous):

Nope. don't treat the statement inside to be seperable.

OpenStudy (anonymous):

Are you familiar with Pacal's triangle?

OpenStudy (zzr0ck3r):

(a+b)*(a+b)*(a+b) start foiling

OpenStudy (anonymous):

foiling?? and no I dont

OpenStudy (zzr0ck3r):

do you know how to do (a+b)*(a+b) = a^2 + 2ab+b^2?

OpenStudy (zzr0ck3r):

so now we have (a+b)*(a^2+2ab+b^2)

OpenStudy (anonymous):

its not 2 it's 3?

OpenStudy (zzr0ck3r):

(a+b)^3 = (a+b)*(a+b)*(a+b) =(a+b)*(a^2 + 2ab+b^2)

OpenStudy (anonymous):

I don't get it. the first part would be a^3+b^3=

OpenStudy (zzr0ck3r):

no (a+b)^3 does not equal a^3+b^3

OpenStudy (zzr0ck3r):

example (2+2)^3 = 4^3 = 64 but 2^3+2^3 = 16

OpenStudy (zzr0ck3r):

what does (a+b)^2 = ?

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

that confused me big time.

OpenStudy (zzr0ck3r):

you have never learned foiling? frst inner outer last?

OpenStudy (anonymous):

y3z3?

OpenStudy (zzr0ck3r):

what is (a+b)^2 ?

OpenStudy (anonymous):

oh 2 a + 2b?

OpenStudy (zzr0ck3r):

no, you might need to back up a bit, google how to foil a binomia

OpenStudy (zzr0ck3r):

binomial*